OpenStudy (anonymous):

how does ln(1/2) = -ln2 ?

myininaya (myininaya):

\[\ln(\frac{a}{b})=\ln(a)-\ln(b)\] \[\ln(1)=0\]

myininaya (myininaya):

\[\ln(\frac{1}{2})=\ln(1)-\ln(2)=0-\ln(2)=-\ln(2)\]

myininaya (myininaya):

or

myininaya (myininaya):

\[\ln(x^r)=r \ln(x)\]

myininaya (myininaya):

\[\ln(\frac{1}{2})=\ln(2^{-1})=(-1)\ln(2)=-\ln(2)\]

OpenStudy (anonymous):

ooo thank you! i really liked ur answers

myininaya (myininaya):

cool!